2x + y = 5; (3, -1)
Please help me with this problem. Thanks

Answer by funmath(2933)

2x + y = 5; (3, -1)
I assume you are being asked to see if (x,y)=(3,-1) is a solution to the equation. To do this, let x=3 and y=-1, simplify and see if both sides equal. If they do, then it is, if they don't then it isn't.
2(3)+(-1)=5
6-1=5
5=5
Yes (3,-1) is a solution to the equation 2x+y=5.
Happy Calculating!!!